# A dynamical systems formulation for inhomogeneous LRS-II spacetimes

### v1

@@ 1 Introduction
In this work, we present a new dynamical system formulation for a class of inhomogeneous spacetimes with a local rotational symmetry, namely the LRS-II. The motivation to consider this class is that it contains the spherically symmetric solutions e.g. LTB that are of physical interest. Our formulation is based on the 1+1+2 covariant decomposition approach [33, 34], which itself is built up on the 1+3 covariant decomposition approach [35, 36]  [^1] That 1+1+2 covariant approach can potentially be helpful in constructing a dynamical system formulation for LRS-II spacetimes is hinted in [37, 38]. However, the authors considered only LRS-II vacuum or electrovacuum solutions for which case it reduces to either static or homogeneous spacetime by virtue of the Birkhoff’s theorem.. To proceed with this approach one needs to first specify a timeline unit vector $u^{\mu}$ and a spacelike unit vector $e^{\mu}$, which we choose to be along the worldline of a comoving observer and the direction of the axis of local rotational symmetry respectively. In the covariant decomposition approaches, all the dynamical quantities have a very clear and coordinate-independent physical interpretation. Physically, our approach involves going into the frame of a comoving observer and following the dynamics from his/her point of view. The effect of spatial inhomogeneity is that not all the comoving observers would experience the same evolutionary phase (characterized by a fixed point), simultaneously, a point we also try to elaborate on in the main text. The novelty of our approach is that we could formulate an autonomous system with only time-dependent evolution equations and algebraic constraints and our dynamical variables are also directly related to covariant geometrical and thermodynamic quantities relevant to a comoving observer.
@@
The paper is organized as follows. In section 2 we present an exposition to LRS-II spacetimes using the 1+1+2 covariant formalism. We dedicate a complete subsection discussing various intricacies presented by the inhomogeneity of the system. In section 3 we construct an autonomous dynamical system for LRS-II using expansion normalized dimensionless dynamical variables. We dedicate an entire subsection to discussing the interpretation of the dynamical structure, something which we deem important here. Sections 4, 5 and 6 present the invariant submanifolds, finite fixed points and asymptotic fixed points in the LRS-II phase space. In section 7 we specialize to the spherically symmetric LTB model and discuss its phase space. We conclude in section 8 summarizing the important and novel aspects of our work with an outlook of possible further research along this line. Finally, appendix A is devoted to exploring a possible different route to constructing an autonomous system for cases with positive 3-curvatures, which proves to be not very fruitful.
@@ 2 Dynamics of LRS-II spacetime with a perfect fluid
where $t,\,r$ are affine parameters along $u^{\mu}$ and $e^{\mu}$ and $k=(-1,0,1)$ denotes open, flat and closed geometry of the 2-sheets respectively. The function $D_{k}(y)$ is
@@
Over the last few years, a new approach, originally introduced by Clarkson [33] has been used extensively in the study of LRS-II spacetimes, both in General Relativity and in $f(R)$ gravity [40, 41, 42, 43] .
The 1+1+2 covariant approach builds on the well-known 1+3 approach to cosmology pioneered by Ehlers [35] and Ellis [36] by further splitting the three spatial degrees of freedom into a part that lies along a spatial vector $e^{a}$ and one perpendicular to $e^{a}$ (referred to as the “sheet”). In this way any spatial tensor can be split into a part along $e^{a}$ and one orthogonal to $e^{a}$.
In general, there is a set of evolution equations containing only ‘dot’-derivatives, a set of propagation equations containing only ‘hat’-derivatives and a set of mixed evolution-propagation equations containing both hat and dot derivatives. The amazing simplicity offered by confining our attention only to perfect fluids is that there we can get rid of the mixed equations. The evolution and the propagation equations can be completely decoupled.
@@
    • $\Sigma\equiv e^{\mu}e^{\nu}\sigma_{\mu\nu}=e^{\mu}e^{\nu}D_{\langle\mu}u_{\nu\rangle}=e^{\mu}e^{\nu}\left(h^{\alpha}_{\,\,\,(\mu}h^{\beta}_{\,\,\,\nu)}-\frac{1}{3}h_{\mu\nu}h^{\alpha\beta}\right)D_{\alpha}u_{\beta}$; projection of the shear tensor $\sigma_{\mu\nu}$ along the spacelike direction $e^{\mu}$.
@@ 4 Invariant submanifolds
    $\frac{\epsilon}{12}\left(3x_{2}-4\pm\sqrt{x_{2}^{2}+24x_{2}}\right),\,\,\frac{\epsilon}{12}\left(5x_{2}-4\pm\sqrt{x_{2}^{2}+24x_{2}}\right)$
@@ 8 Conclusion
We identify several invariant submanifolds and fixed points in sections 4, 5 and 6. We found that homogeneous solutions contain an invariant submanifold whose stability depends only on the quantity $\frac{\mu}{\theta^{2}}$ (or equivalently $\frac{{}^{3}R}{\theta^{2}}$). For LTB, one can qualitatively say that a large matter density hinders the homogenization of an almost-FLRW expanding LTB whereas assists in the homogenization of an almost-FLRW contracting LTB. Interestingly, we find that spatially flat solutions in general do not constitute an invariant submanifold in the LRS-II phase space because of a non-vanishing shear. Finding a solution corresponding to vacuum fixed points becomes ambiguous due to not having a preferred choice for $u^{\mu}$ (i.e. a preferred frame), something which is a precondition for the applicability of the 1+3 formalism. Special care is taken to interpret the solutions corresponding to the vacuum fixed points. We find a spatially flat matter-dominated FLRW solution as a saddle fixed point, i.e. an intermediate epoch of evolution, as one might expect in the presence of nonrelativistic matter. We also recovered a future attractor which a comoving observer would identify with Milne, consistent with earlier results that expanding LTB solutions without the cosmological constant tends to Milne [46, 47].

### v2

@@ 1 Introduction
In this work, we present a new dynamical system formulation for a class of inhomogeneous spacetimes with a local rotational symmetry, namely the LRS-II. The motivation to consider this class is that it contains the spherically symmetric solutions e.g. LTB that are of physical interest. Our formulation is based on the 1+1+2 covariant decomposition approach [33, 34], which itself is built up on the 1+3 covariant decomposition approach [35, 36]  [^1] That 1+1+2 covariant approach can potentially be helpful in constructing a dynamical system formulation for LRS-II spacetimes is hinted in [37, 38]. However, the authors considered only LRS-II vacuum or electrovacuum solutions for which case it reduces to either static or homogeneous spacetime by virtue of the Birkhoff’s theorem.. To proceed with this approach one needs to first specify a timeline unit 4-vector $u^{\mu}$ and a spacelike unit 4-vector $e^{\mu}$, which we choose to be along the worldline of a comoving observer and the direction of the axis of local rotational symmetry respectively. In the covariant decomposition approaches, all the dynamical quantities have a very clear and coordinate-independent physical interpretation. Physically, our approach involves going into the frame of a comoving observer and following the dynamics from his/her point of view. The effect of spatial inhomogeneity is that not all the comoving observers would experience the same evolutionary phase (characterized by a fixed point), simultaneously, a point we also try to elaborate on in the main text. The novelty of our approach is that we could formulate an autonomous system with only time-dependent evolution equations and algebraic constraints and our dynamical variables are also directly related to covariant geometrical and thermodynamic quantities relevant to a comoving observer.
@@
The paper is organized as follows. In section 3 we present an exposition to LRS-II spacetimes using the 1+1+2 covariant formalism. We dedicate a complete subsection discussing various intricacies presented by the inhomogeneity of the system. In section 4 we construct an autonomous dynamical system for LRS-II using expansion normalized dimensionless dynamical variables. We dedicate an entire subsection to discussing the interpretation of the dynamical structure, something which we deem important here. Sections 5, 6 and 7 present the invariant submanifolds, finite fixed points and asymptotic fixed points in the LRS-II phase space. In section 8 we specialize to the spherically symmetric LTB model and discuss its phase space. We conclude in section 9 summarizing the important and novel aspects of our work with an outlook of possible further research along this line. Finally, appendix A is devoted to exploring a possible different route to constructing an autonomous system for cases with positive 3-curvatures, which proves to be not very fruitful.
2 Some basics about covariant splitting
The well-known 1+3 covariant approach to cosmology pioneered by Ehlers [35] and Ellis [36] starts by splitting all the covariant geometric and thermodynamic quantities into a part that lies along a preferred timelike unit 4-vector $u^{\mu}$ ($u^{\mu}u_{\mu}=-1$) and one perpendicular to it. The specific choice of a 4-vector essentially chooses a “frame”. The metric tensor $g_{\mu\nu}$ can be decomposed as
$g_{\mu\nu}=U_{\mu\nu}+h_{\mu\nu},$ (2.1)
where the tensor $U_{\mu\nu}=-u_{\mu}u_{\nu}$ projects along $u^{\mu}$ and the tensor $h_{\mu\nu}$ projects on the 3-surface orthogonal to $u^{\mu}$ ($h_{\mu\nu}u^{\nu}=0$). A 4-scalar quantity cannot be further decomposed. A 4-vector $V^{\mu}$ can be irreducibly decomposed into a scalar and a 3-vector using Eq.(2.1):
$V^{\mu}=-\mathcal{V}u^{\mu}+\mathcal{V}^{\mu}\;,$ (2.2)
where the scalar $\mathcal{V}=V^{\mu}u_{\mu}$ is the projection of $V^{\mu}$ along the 4-vector $u^{\mu}$ and the 3-vector $\mathcal{V}^{\mu}=h^{\mu}_{\alpha}V^{\alpha}$ is the projection of $V^{\mu}$ on the 3-surface orthogonal to $u^{\mu}$. In the language of covariant 1+3 splitting, we can say $\mathcal{V}^{\mu}$ is a 3-vector because it completely lies on the 3-surface orthogonal to $V^{\mu}$ ($\mathcal{V}^{\mu}u_{\mu}=0$).
A 4-tensor can be irreducibly decomposed into a trace part, which is a scalar, and a tracefree part. The tracefree part can be further decomposed using Eq.(2.1). For example, a tracefree 2nd rank 4-tensor $T_{\mu\nu}$ can be irreducibly decomposed into a scalar, 3-vectors and a projected tracefree 3-tensor using Eq.(2.1) as follows:
$T_{\mu\nu}=\mathcal{T}\left(u_{\mu}u_{\nu}+\frac{1}{3}h_{\mu\nu}\right)+\mathcal{T}_{1\mu}u_{\nu}+\mathcal{T}_{2\nu}u_{\mu}+\mathcal{T}_{\mu\nu},$ (2.3)
where we have defined
$\mathcal{T}\equiv T^{\alpha\beta}u_{\alpha}u_{\beta}=T^{\alpha\beta}h_{\alpha\beta}\;,$ (2.4a)
$\mathcal{T}_{1\mu}\equiv-T^{\alpha\beta}h_{\alpha\mu}u_{\beta},\quad\mathcal{T}_{2\mu}\equiv-T^{\alpha\beta}u_{\alpha}h_{\beta\mu}\;,$ (2.4b)
$\mathcal{T}_{\mu\nu}\equiv\left(h_{\mu\alpha}h_{\nu\beta}-\frac{1}{3}h_{\mu\nu}h_{\alpha\beta}\right)T^{\alpha\beta}\;.$ (2.4c)
In the language of the 1+3 covariant splitting, we can say $\mathcal{T}_{1\mu},\,\mathcal{T}_{2\mu}$ and $\mathcal{T}_{\mu\nu}$ are 3-vectors and 3-tensors, respectively, because they lie completely on the 3-surface orthogonal to $u^{\mu}$ ($\mathcal{T}_{1\mu}u^{\mu}=0=\mathcal{T}_{2\mu}u^{\mu}$, $\mathcal{T}_{\mu\nu}u^{\mu}=0$). Also, note that the 3-tensor $\mathcal{T}_{\mu\nu}$ is tracefree ($h_{\mu\nu}\mathcal{T}^{\mu\nu}=0$).
The covariant derivative of a 4-vector and a 2nd rank 4-tensor can be decomposed respectively as
$\nabla_{\mu}V^{\alpha}=-\dot{V}^{\alpha}u_{\mu}+D_{\mu}V^{\alpha}-u^{\alpha}u_{\gamma}h^{\nu}_{\,\,\,\mu}\nabla_{\nu}V^{\gamma}\,,$ (2.5a)
$\nabla_{\mu}T^{\alpha}_{\,\,\,\beta}=-\dot{T}^{\alpha}_{\,\,\,\beta}u_{\mu}+D_{\mu}T^{\alpha}_{\,\,\,\beta}+h^{\nu}_{\,\,\,\mu}u^{\alpha}u^{\sigma}u_{\gamma}u_{\beta}\nabla_{\nu}T^{\gamma}_{\,\,\,\sigma}-(u^{\alpha}u_{\gamma}h^{\sigma}_{\,\,\,\beta}+u^{\sigma}u_{\beta}h^{\alpha}_{\,\,\,\gamma})h^{\nu}_{\,\,\,\mu}\nabla_{\nu}T^{\gamma}_{\,\,\,\sigma}\;,$
where we have defined the covariant derivative along the timelike 4-vector $u^{\mu}$ and the fully orthogonally projected covariant derivative on the 3-surface respectively as
$\dot{V}^{\alpha}=u^{\mu}\nabla_{\mu}V^{\alpha}\,,$ $\dot{T}^{\alpha}_{\,\,\,\beta}=u^{\mu}\nabla_{\mu}T^{\alpha}_{\,\,\,\beta}\,,$ (2.6a)
$D_{\mu}V^{\alpha}=h^{\nu}_{\,\,\,\mu}h^{\alpha}_{\,\,\,\gamma}\nabla_{\nu}V^{\gamma}\,,$ $D_{\mu}T^{\alpha}_{\,\,\,\beta}=h^{\nu}_{\,\,\,\mu}h^{\alpha}_{\,\,\,\gamma}h^{\sigma}_{\,\,\,\beta}\nabla_{\nu}T^{\gamma}_{\,\,\,\sigma}\,.$ (2.6b)
It is straightforward to generalize the above definitions for higher-rank 4-tensors.
The 1+1+2 covariant approach [33, 34] builds on the 1+3 approach by further splitting the three spatial degrees of freedom into a part that lies along a preferred spacelike 4-vector $e^{\mu}$ and one perpendicular to $e^{\mu}$ (referred to as the “sheet”). The vector $e^{\mu}$ is taken to be normalized and it lies completely on the 3-surface orthogonal to $u^{\mu}$
$e^{\mu}e_{\mu}=1,\quad u^{\mu}e_{\mu}=0.$ (2.7)
In the language of 1+3 covariant splitting, we can say $e^{\mu}$ is a 3-vector. With respect to $e^{\mu}$, the 3-tensor $h_{\mu\nu}$ is further decomposed as
$h_{\mu\nu}=E_{\mu\nu}+N_{\mu\nu},$ (2.8)
where the 3-tensor $E_{\mu\nu}=e_{\mu}e_{\nu}$ projects along $e^{\mu}$ and the 3-tensor $N_{\mu\nu}$ projects on the 2-sheet orthogonal to both $u^{\mu}$ and $e^{\mu}$ ($N_{\mu\nu}u^{\mu}=0=N_{\mu\nu}e^{\mu}$). Using Eq. (2.8) a 3-vector $\mathcal{V}^{\mu}$ can be irreducibly decomposed into a scalar and a 2-vector
$\mathcal{V}^{\mu}=\mathbb{V}e^{\mu}+\mathbb{V}^{\mu}\;,$ (2.9)
where the scalar $\mathbb{V}=\mathcal{V}^{\mu}e_{\mu}$ is the projection of the 3-vector $\mathcal{V}^{\mu}$ along the 3-vector $e^{\mu}$ and the 2-vector $\mathbb{V}^{\mu}=N^{\mu\alpha}\mathcal{V}_{\alpha}$ is the projection of the 3-vector $\mathcal{V}^{\mu}$ on the 2-dimensional sheet orthogonal to $e^{\mu}$. We say $\mathbb{V}^{\mu}$ is a 2-vector because it completely lies on the 2-sheet orthogonal to both $u^{\mu}$ and $e^{\mu}$ ($\mathbb{V}^{\mu}u_{\mu}=0=\mathbb{V}^{\nu}e_{\nu}$).
Similarly, a tracefree 3-tensor can be further decomposed using Eq. (2.8). For example, a tracefree 2nd rank 3-tensor $\mathcal{T}_{\mu\nu}$ can be irreducibly decomposed into a scalar, 2-vectors and a projected tracefree 2-tensor using Eq. (2.8) as follows:
$\mathcal{T}_{\mu\nu}=\mathbb{T}\left(e_{\mu}e_{\nu}-\frac{1}{2}N_{\mu\nu}\right)+\mathbb{T}_{1\mu}e_{\nu}+\mathbb{T}_{2\nu}e_{\mu}+\mathbb{T}_{\mu\nu},$ (2.10)
where we have defined
$\mathbb{T}\equiv\mathcal{T}^{\alpha\beta}e_{\alpha}e_{\beta}=-\mathcal{T}^{\alpha\beta}N_{\alpha\beta},$ (2.11a)
$\mathbb{T}_{1\mu}\equiv\mathcal{T}^{\alpha\beta}N_{\alpha\mu}e_{\beta},\quad\mathbb{T}_{2\mu}\equiv\mathcal{T}^{\alpha\beta}e_{\alpha}N_{\beta\mu}$ (2.11b)
$\mathbb{T}_{\mu\nu}\equiv\left(N_{\mu\alpha}N_{\nu\beta}-\frac{1}{2}N_{\mu\nu}N_{\alpha\beta}\right)\mathcal{T}^{\alpha\beta}$ (2.11c)
$\mathbb{T}_{1\mu},\,\mathbb{T}_{2\mu}$ and $\mathbb{T}_{\mu\nu}$ are 2-vector and 2-tensor, respectively, because they lie completely on the 2-sheet orthogonal to $u^{\mu}$ and $e^{\mu}$ ($\mathbb{T}_{1\mu}u^{\mu}=\mathbb{T}_{1\mu}e^{\mu}=0=\mathbb{T}_{2\mu}e^{\mu}=\mathbb{T}_{2\mu}u^{\mu}$, $\mathbb{T}_{\mu\nu}u^{\mu}=0=\mathbb{T}_{\mu\nu}e^{\mu}$). Also, note that the 2-tensor $\mathbb{T}_{\mu\nu}$ is tracefree ($N_{\mu\nu}\mathbb{T}^{\mu\nu}=0$).
The fully orthogonally projected covariant derivative of a 3-vector and a 2nd rank 3-tensor can be decomposed respectively as
$D_{\mu}\mathcal{V}^{\alpha}=\hat{\mathcal{V}}^{\alpha}e_{\mu}+d_{\mu}\mathcal{V}^{\alpha}+e^{\alpha}e_{\gamma}N^{\nu}_{\,\,\,\mu}D_{\nu}V^{\gamma}\,,$ (2.12a)
$D_{\mu}\mathcal{T}^{\alpha}_{\,\,\,\beta}=\hat{\mathcal{T}}^{\alpha}_{\,\,\,\beta}e_{\mu}+d_{\mu}\mathcal{T}^{\alpha}_{\,\,\,\beta}+N^{\nu}_{\,\,\,\mu}e^{\alpha}e^{\sigma}e_{\gamma}e_{\beta}D_{\nu}\mathcal{T}^{\gamma}_{\,\,\,\sigma}+(e^{\alpha}e_{\gamma}N^{\sigma}_{\,\,\,\beta}+e^{\sigma}e_{\beta}N^{\alpha}_{\,\,\,\gamma})N^{\nu}_{\,\,\,\mu}D_{\nu}\mathcal{T}^{\gamma}_{\,\,\,\sigma}\,,$
where we have defined the covariant derivative along the spacelike 3-vector $e^{\mu}$ and the fully orthogonally projected covariant derivative on the 2-sheet respectively as:
$\hat{\mathcal{V}}^{\alpha}=e^{\mu}D_{\mu}\mathcal{V}^{\alpha}\,,$ $\hat{\mathcal{T}}^{\alpha}_{\,\,\,\beta}=e^{\mu}D_{\mu}\mathcal{T}^{\alpha}_{\,\,\,\beta}\,,$ (2.13a)
$d_{\mu}\mathcal{V}^{\alpha}=N^{\nu}_{\,\,\,\mu}N^{\alpha}_{\,\,\,\gamma}D_{\nu}\mathcal{V}^{\gamma}\,,$ $d_{\mu}\mathcal{T}^{\alpha}_{\,\,\,\beta}=N^{\nu}_{\,\,\,\mu}N^{\alpha}_{\,\,\,\gamma}N^{\sigma}_{\,\,\,\beta}D_{\nu}\mathcal{T}^{\gamma}_{\,\,\,\sigma}\,.$ (2.13b)
It is straightforward to generalize the above definitions for higher-rank 3-tensors.
Eqs. (2.1) and (2.8) can be written in a combined form as
$g_{\mu\nu}=U_{\mu\nu}+E_{\mu\nu}+N_{\mu\nu}=-u_{\mu}u_{\nu}+e_{\mu}e_{\nu}+N_{\mu\nu},$ (2.14)
which can be used to decompose all the covariant physical quantities into scalars, 2-vectors and 2-tensors, with the later two lying completely on the 2-dimensional sheet orthogonal to $u^{\mu},\,e^{\mu}$. The beautiful aspect of the 1+1+2 decomposition formulation is that, for spacetimes with a local rotational symmetry, if the spatial 2-vector $e^{\mu}$ is taken at each spatial point to be along the axis of the local rotational symmetry, all the 2-vectors and 2-tensors vanish, leaving us only with a set of scalars that characterize the dynamics.
When the metric is diagonal and the temporal and the radial coordinates are affine parameters along $u^{\mu}$ and $e^{\mu}$ respectively, then the explicit form of these vectors are
$u^{\mu}=\left(\frac{1}{\sqrt{-g_{00}}},0,0,0\right)\,,\quad e^{\mu}=\left(0,\frac{1}{\sqrt{g_{rr}}},0,0\right)\,.$ (2.15)
The covariant derivatives along these two vectors for a scalar quantity are related to its ordinary time and radial derivatives as follows:
$\dot{\psi}=u^{\mu}\nabla_{\mu}\psi=u^{\mu}\partial_{\mu}\psi=\frac{1}{\sqrt{-g_{00}}}\partial_{t}\psi\,,$ (2.16a)
$\hat{\psi}=e^{\mu}D_{\mu}\psi=e^{\mu}h^{\nu}_{\mu}\nabla_{\nu}\psi=e^{\mu}\partial_{\mu}\psi=\frac{1}{\sqrt{g_{rr}}}\partial_{r}\psi\,.$ (2.16b)
The above formulation is completely covariant, i.e. coordinate-independent. When a particular coordinate system is chosen and the metric is specified, different covariant quantities can be calculated explicitly in terms of the metric functions.
@@ 3 Dynamics of LRS-II spacetime with a perfect fluid
where $t,\,r$ are affine parameters along the preferred timeline 4-vector $u^{\mu}$ and the preferred spacelike 4-vector $e^{\mu}$ and $k=(-1,0,1)$ denotes open, flat and closed geometry of the 2-sheets respectively. The function $D_{k}(y)$ is
@@
Over the last few years, the covariant 1+1+2 splitting approach, originally introduced by Clarkson in Ref.[33] to study perturbations around the Schwarzschild background, has been used extensively in the study of LRS-II spacetimes, both in General Relativity and in $f(R)$ gravity [40, 41, 42, 43]. In general, for LRS-II spacetimes, there is a set of evolution equations containing only ‘dot’-derivatives, a set of propagation equations containing only ‘hat’-derivatives and a set of mixed evolution-propagation equations containing both hat and dot derivatives. The amazing simplicity offered by confining our attention only to perfect fluids is that there we can get rid of the mixed equations. The evolution and the propagation equations can be completely decoupled.
@@
    • $\Sigma\equiv e^{\mu}e^{\nu}\sigma_{\mu\nu}=e^{\mu}e^{\nu}D_{\langle\mu}u_{\nu\rangle}=e^{\mu}e^{\nu}\left(h^{\alpha}_{\,\,\,(\mu}h^{\beta}_{\,\,\,\nu)}-\frac{1}{3}h_{\mu\nu}h^{\alpha\beta}\right)D_{\alpha}u_{\beta}$; projection of the shear tensor $\sigma_{\mu\nu}=e^{\mu}e^{\nu}D_{\langle\mu}u_{\nu\rangle}$ along the spacelike direction $e^{\mu}$ 222delimited-⟨$\langle$..$\rangle$ … ⟩ in the indices denotes symmetrization of indices..
@@ 5 Invariant submanifolds
    $\frac{\epsilon}{12}\left(3x_{2}-4\pm\sqrt{x_{2}^{2}+24x_{2}}\right),\,\,\frac{\epsilon}{12}\left(5x_{2}-4\pm\sqrt{x_{2}^{2}+24x_{2}}\right)\;.$
@@ 9 Conclusion
We identify several invariant submanifolds and fixed points in sections 5, 6 and 7. We found that homogeneous solutions contain an invariant submanifold whose stability depends only on the quantity $\frac{\mu}{\theta^{2}}$ (or equivalently $\frac{{}^{3}R}{\theta^{2}}$). For LTB, one can qualitatively say that a large matter density hinders the homogenization of an almost-FLRW expanding LTB whereas assists in the homogenization of an almost-FLRW contracting LTB. Interestingly, we find that spatially flat solutions in general do not constitute an invariant submanifold in the LRS-II phase space because of a non-vanishing shear. Finding a solution corresponding to vacuum fixed points becomes ambiguous due to not having a preferred choice for $u^{\mu}$ (i.e. a preferred frame), something which is a precondition for the applicability of the 1+3 formalism. Special care is taken to interpret the solutions corresponding to the vacuum fixed points. We find a spatially flat matter-dominated FLRW solution as a saddle fixed point, i.e. an intermediate epoch of evolution, as one might expect in the presence of nonrelativistic matter. We also recovered a future attractor which a comoving observer would identify with Milne, consistent with earlier results that expanding LTB solutions without the cosmological constant tends to Milne [46, 47].